Convert equation to standard form for a circle if needed then solve center and radius for four problem
1) x^2+y^2+6y=-50-14x
2) x^2+y^2+2x-15=0
3) x^2+y^2-3x+8y=20
4) x^2+y^2+2x+4y=20

Answer by lwsshak3(11628)

Convert equation to standard form for a circle if needed then solve center and radius for four problem
standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=coordinates of center, r=radius
1) x^2+y^2+6y=-50-14x
x^2+14x+y^2+6y=-50
complete the square:
(x^2+14x+49)+(y^2+6y+9)=-50+49+9
(x+7)^2+(y+3)^2=8
center: (-7,-3), radius=√8
..
2) x^2+y^2+2x-15=0
complete the square
(x^2+2x+1)+y^2=15+1
(x+1)^2+y^2=16
center:(-1,0), radius=4
..
let you do the rest
3) x^2+y^2-3x+8y=20
4) x^2+y^2+2x+4y=20